A 232Th atom (Z=90, Atomic mass = 232.038054 amu) consists of protons (Mass = 1.00728 amu), neutrons (Mass = 1.00867 amu) and electrons (Mass = 0.000549 amu). Calculate the binding energy per nucleon.
Answer #1

Nuclear mass = 232.03854amu
No. of nucleon = 232
proton = A = 90 , Neutron = A-Z = 232-90=142
nucleon mass = no. of proton * mass of proton + no. of neutron *mass of neutron
so 90*1.00728+142*1.00867 = 90.6552+143.2311=233.88634amu
Binding Energy is
BE =(mass of nucleon - nuclear mass ) 931.5 MeV
=(233.88634-232.038054)*931.5 = 1.84825*931.5 = 1721.678409MeV
BE/no. of nucleon = 1721.678409/232=7.42102 MeV/amu

if you need in joule then

(7.421MeV)(1.6x10^-13 J/MeV) = 1.1873x10^-12J